Carpet sells for $3 per square foot and will cost you $546 to recarpet your room. If your room is 14 feet long, how many feet wide is it?

Found 2 solutions by leducanh1992, rfer:
Answer by leducanh1992(9)

The are of the room is 546/3 = 182 sqft. The width of the room is Area/Length = 182/14 = 13 ft.

Answer by rfer(16322)

546/3=182 sqft
182/14=13 ft